OpenStudy (megannicole51):

noel currently has a balance of 2300.37 in an account he has held for 34 years. He opened the account with an initial deposit of 874.00...what is the simple interest rate on the account?

OpenStudy (megannicole51):

thank you!

OpenStudy (anonymous):

Sorry. There was a slight mistake in my previous answer. Here is the correct answer: We use the simple interest formula: Total Interest = Initial Deposit * Rate of Increase * Time. Total Interest = Current Balance - Initial Deposit Total Interest = 2300.37 - 874 = 1426.37 Plug the numbers into the equation. 1426.37 = 874 * Rate of Increase * 34 1426.37 = 29716 * Rate of Increase Rate of Increase = Approximately 0.048 = 4.8% The answer is 4.8% .

OpenStudy (anonymous):

Let me know if you need further explanation.

OpenStudy (anonymous):

You can also check your answer by doing this: 1.) 0.048 * 874 = 41.952 (this is the amount gained each year and since there are 34 years, we move on to step 2. . .) 2.) 41.952 * 34 = 1426.368 (Notice this as the total interest) 3.) 1426.368 + 874 = 2300.368 (Our total amount -- Which is pretty close!)